# Consistent Validation for Predictive Methods in Spatial Settings

### v2

@@ 6.5 Air Temperature Prediction with Real Response
We finally consider a case with real data and a real task. Although we cannot access ground truth, we show that the holdout, 1NN, and our SNN give very different estimates of test risk and can differ in model selection. Given all the previous results, we suggest using SNN. See Section E.4 for full details and also a grid-prediction task where all estimators are in agreement.
@@
Data and Models Fit. The data, test task and models considered are the same as in Section 6.2, but the actual response values are used for training and validation.
@@
Acknowledgements
@@
This work was supported in part by an NSF CAREER Award and the Office of Naval Research under grant N00014-20-1-2023 (MURI ML-SCOPE).
@@
Appendix A Broader Impact Statement
This paper identifies potential failure modes of existing validation techniques in a spatial setting and suggests practical improvements, with supporting theory and empirics. We hope that our work can play a part in improving the reliability of prediction assessment and thereby help improve the credibility of scientific analyses.
Appendix B Extended Related Work
@@
B.1 Validation versus Cross-Validation
@@
While we expect our work to have implications for cross validation (CV), we focus on validation here since (1) we do not assume training data exists or is easy to change (for example in the case of a large physical model) and (2) CV presents additional subtle challenges. The holdout is broadly applicable to any predictive method (whether data-driven, physical, or a combination thereof). However, in the cases when data is scarce and a data-driven predictive method is used, cross-validation is commonly believed to make more efficient use of the data. CV is widely used in spatial analyses; among many examples are Wang et al. [2022], Valavi et al. [2021], Kianian et al. [2021]. However, the interpretation of the error estimates given by cross-validation is subtle even in the classical i.i.d. setting [Bates et al., 2023]. We focus on the validation setting in this work due to its broad applicability and clear interpretation, though extensions to cross-validation, as well as a theoretical understanding of how the resulting error estimate relate to predictive performance is a promising direction for future work.
B.2 Covariate Shift
@@
In covariate shift, it is generally assumed that $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}P^{\text{val}}$, $S^{\textup{{test}}}_{m}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}Q$ and $U^{\ell}_{i}|S^{\ell}_{i}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}\mu$ for $\ell\in\{\text{val},\text{test}\}$. Moreover, it is typically assumed the density ratio $\frac{dQ}{dP}$ exists and is bounded, although there is recent work relaxing this assumption in the context of non-parametric regression [Kpotufe and Martinet, 2021, Pathak et al., 2022]. Mean estimation seeks to estimate $\mathbb{E}[U^{\text{test}}_{1}]$. Taking $(U^{\text{test}}_{m}|S^{\textup{{test}}}_{m})=\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))$, this is the same task we consider, but with the (stronger) assumption that covariates are independent and identically distributed with a bounded density ratio between the test and validation distributions. This assumptions is not appropriate for validation with many spatial datasets, with particularly simple examples being when the task of interest requires prediction at a single spatial location, or on a regular grid. Our assumptions are essentially a conditional formulation of mean estimation under covariate shift. Many methods proposed for addressing mean estimation under covariate shift, including all the approaches we describe below, are based on re-weighting validation points, then applying the holdout approach described earlier with these weights.
@@
The kernel mean matching algorithm provides a solution under standard covariate shift assumptions with a bounded density ratio and assuming the average loss as a function of space $e_{h}$ lives in a reproducing kernel Hilbert space (RKHS) and has small norm [Gretton et al., 2009]. Yu and Szepesvári [2012] provide a finite sample bound for this method
Appendix C Error Estimation for Model Selection
@@
We are not able to formally characterize how SNN performs in model selection. But we provide rigorous results on rates of convergence of SNN that give suggestive guidance. In particular, (a) we first consider the case of a grid prediction task, where test sites are arranged on a grid. In this case, we are able to give a rate of convergence for both $1$NN and SNN (Corollaries D.16 and D.17). If the validation data are i.i.d. our rate is faster than the minimax optimal convergence rate for predicting a Lipschitz function in the presence of additive, homoskedastic noise. Our result shows the same rate of convergence as Portier et al. [2023, Prop. 4], who considered test data that was i.i.d. instead of on a grid. Our result is suggestive that both nearest neighbor methods may perform well at model selection for validating maps (grid prediction).
(b) Second, we provide finite-sample bounds and asymptotic characterizations for general (non-grid) $Q^{\textup{{test}}}$ (Corollary D.14). In this case, when validation data is i.i.d., up to logarithmic factors, we show that SNN converges at the optimal rate of convergence for Lipschitz functions (Corollary D.18). Since in this case, the SNN convergence rate is not strictly faster than the training convergence rate, it might be difficult to select between statistical methods that converge at the optimal regression rate. But we would still expect to be able to select between (i) a statistical method that converges at the optimal rate and (ii) one that does not (for example a misspecified parametric model). Since the holdout and 1NN may not even be consistent, we could not rely on them to select the better model even in this latter, easier case.
Appendix D Proofs of Claims
We now present results and proofs not included in the main text. We essentially follow the order of results in the main text. Section Section D.1 gives sufficient conditions for Assumption 2.5 to hold for a homoskedastic, additive noise model and squared loss for responses taking values in $[0,1]$, mentioned in Section 2.2. Section D.2 focuses on the proof of Proposition 3.3, first recalling several properties of covering numbers that will be used in the result. In Section D.3 we restate and prove our Propositions 4.1, 4.3 and 4.4, which show limitations of existing validation methods in the spatial setting we study. In Section D.4 we prove a general result upper bounding the error in estimating the risk using $k$-nearest neighbors, as well as an upper bound on the error for $k^{\star}_{T_{2}}$. Section D.5 establishes the consistency of spatial nearest neighbors. Section D.6 discusses issues related to model selection and proves rates of convergence for spatial nearest neighbors.
D.1 Lipschitz Constant for Lipschitz Response and Predictive Method
@@
D.2 Proof that Independent and Identically Distributed Data Implies Infill Asymptotic with High Probability
@@
Our proof is similar to earlier proofs in Reznikov and Saff [2015, Section 5.2] and essentially follows the stackoverflow response (2021) [https://mathoverflow.net/users/36721/iosif pinelis] keeping track of numerical constants. Essentially, the idea is that were the fill distance to be large, there must be a ball of large radius that doesn’t contain any points in the sample of location. But because the probability of a point in the sample falling in any ball is bounded below in terms of the volume of the ball, this must be improbable as the sample size grows.
D.2.1 Preliminary Definitions and Lemmas Related to Nets and Covering Number
@@
Definition D.2 (Net, Covering Number).
@@
Lemma D.4 (Covering number of Unit Cube).
@@
By the upper bound in Wainwright [2019, Lemma 5.7],
@@ D.2.2 Main Proof
Returning to Equation 16 and taking a union over all elements in the net:
@@
where $B_{d}=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2}+1)}$ is the volume of the $d$-dimensional unit sphere. Equation 19 uses Equation 17. We now upper bound the terms in Equation 20.
Applying, Propositions D.3 and D.4
@@
D.3 Proof of Inconsistency of Existing Methods
@@ D.3.1 Inconsistency of the Holdout
Proposition D.5 (Counterexample to the consistency of holdout).
@@
Combining Equation 26 and Equation 27, for all $N\geq N_{0}$ with probability at least $1/2$
@@
D.3.2 Inconsistency of $1$-nearest neighbor Estimator
We now turn to $1$-nearest neighbor risk estimation and restate and prove Proposition 4.3: See 4.3 We again actually prove a strong result
Proposition D.7.
@@ D.3.3 Preliminary Result
Take $t=\frac{A^{2}-C}{\delta}$ which is greater than 00 because $\delta\in(1-\frac{C}{A^{2}},1)$. Then Equation 32 becomes,
@@
D.3.4 Proof of Inconsistency of $1$-nearest neighbor
We return to our proof of Proposition 4.3. The idea will be to consider a point prediction task and then apply Proposition D.8 to show that with some fixed probability, the $1$-nearest neighbor estimator is a fixed distance away from the test risk, even as $N^{\textup{{val}}}$ increases.
@@ D.3.4 Proof of Inconsistency of $1$-nearest neighbor
We now apply Proposition D.8 with $U=|R_{Q^{\textup{{test}}}}(h)-(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)|$ to conclude that for $\delta\in(1-\frac{VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}}{\Delta^{2}},1)$ with probability at least $1-\delta$
@@
As neither $\delta$ nor the right hand side of Equation 42 depend on $N^{\textup{{val}}}$, $1$-nearest neighbor is not consistent under infill asymptotics. ∎
Proposition D.9.
Consider $\mathcal{Y}=\{0,1\}$ and $\ell(a,b)=\begin{cases}0&a=b\\ 1&\text{otherwise}\end{cases}$, $\mathcal{S}=[0,1]^{d}$ and $S^{\textup{{val}}}_{m}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}\mu$, where $\mu$ is any measure with density with respect to Lebesgue measure. Fix any $S\in\mathcal{S}$ and any predictive method $h$ and let $Q^{\textup{{test}}}=\delta_{S}$. Then,
@@
Because $S^{\textup{{val}}}_{m}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}\mu$, and $\mu$ has Lebesgue density, the nearest neighbor to $S$ is almost surely unique. This implies that $\hat{R}_{\textup{{NN}},1}(h)\in\{0,1\}$. Therefore,
@@
D.3.5 Inconsistency of Nearest Neighbors with Number of Neighbors Depending On Number of Validation Points

### v3

@@ 6.5 Temperature, Real Response
We finally consider a case with real data and a real task. Although we cannot access ground truth, we show that the holdout, 1NN, and our SNN give very different estimates of test risk and can differ in model selection. Given all the previous results, we suggest using SNN. See section D.4 for full details and a grid-prediction task where all estimators are in agreement.
@@
Data and Models. The data, models and test task are the same as in section 6.2, but the actual response values are used for training and validation.
@@
7 DISCUSSION
In this paper, we precisely formulate the problem of validation in the spatial setting; we provide a new check for validating spatial predictions (consistency under infill asymptotics); we demonstrate empirically and theoretically that common existing methods don’t pass this check; and we show by construction that it is possible to pass the check with a practical algorithm (our SNN algorithm).
Validation of Spatio-temporal Forecasts. Many prediction tasks have a temporal as well as a spatial component. Our method can be applied when treating time as part of a “location,” but infill asymptotics would then imply we expect access to arbitrarily dense observations in time. We believe it is more typical in temporal prediction tasks to accrue data forward in time and to be interested in prediction forward in time. In the spirit of the present work, one might formalize the intuition that past forecasting performance is informative about future performance as a stationarity assumption.
Validation for Extrapolation. Even in spatial tasks, practitioners often want to extrapolate to unobserved parts of the space. Indeed, blocked spatial validation is often motivated by this goal. We expect consistent risk estimation in this setting to require additional assumptions about the data-generating process (beyond our smoothness assumption, assumption 2.5). We also expect that blocked spatial validation may still be practically useful in extrapolation settings, as discussed in Roberts et al. (2017). Consistent estimation of predictive error in extrapolation settings (either forward in time or across space) seems a challenging, but important, area for future research.
Cross-validation. We have focused on validation of a fixed predictive method. When statistical or machine learning methods are used for prediction, and data is scarce, cross-validation is frequently used. However, recent research has recognized that cross-validation actually incorporates an interesting, non-trivial additional layer of complexity beyond validation (Bates et al., 2023). Indeed, this research has shown that even in the classic independent and identically distributed setting, cross-validation is not estimating the same quantity as validation. So we expect a full understanding of cross-validation in the spatial setting to similarly require substantial additional work beyond understanding validation.
Acknowledgments
@@
This work was supported in part by an NSF CAREER Award, the Office of Naval Research under grant N00014-20-1-2023 (MURI ML-SCOPE), Generali, and a Microsoft Trustworthy AI Grant.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes. Assumptions needed for our algorithm to be consistent are stated in the main text and are linked in the following result statement: corollary 5.2.
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. We derive the computational complexity of our algorithm in section D.2.1, see proposition D.1.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes, code is available at: https://github.com/DavidRBurt/Consistent-Spatial-Validation. Dependencies can be found in the setup file, and instructions to run the code are in the readme file.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Yes, we state all assumptions in theorem statements.
    2. (b)
      Complete proofs of all theoretical results. Yes, proofs are in appendix C.
    3. (c)
      Clear explanations of any assumptions. Yes, we have provided intuition for, and descriptions of, assumptions 2.2, 2.1, 2.4 and 2.5 around the statement of these assumptions.
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes, see prior answer regarding code. Experimental details are included in appendix D.
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes, all training details are provided in appendix D.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Yes, we provide detailed descriptions of the types of randomness considered in each experiment (e.g. bootstrapping versus generating completely independent synthetic experiments) in appendix D, as well as the method and limitations for estimating the “ground truth” risk.
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes, we describe computing infrastructure in section D.2.2.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Yes, we provide citations for datasets in sections D.4.1, D.5.1 and D.6.1. We include citations for libraries used as well.
    2. (b)
      The license information of the assets, if applicable. Yes, we describe license information of datasets in sections D.4.1, D.5.1 and D.6.1.
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. The only new asset is code. See earlier responses.
    4. (d)
      Information about consent from data providers/curators. Not applicable. All data used is permitted by terms of licensing. Most is open data (product of government agency).
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
@@
Appendix A EXTENDED RELATED WORK
@@
A.1 Validation versus Cross-Validation
@@
While we expect our work to have implications for cross validation (CV), we focus on validation here since (1) we do not assume training data exists or is easy to change (for example in the case of a large physical model) and (2) CV presents additional subtle challenges. The holdout is broadly applicable to any predictive method (whether data-driven, physical, or a combination thereof). However, in the cases when data is scarce and a data-driven predictive method is used, cross-validation is commonly believed to make more efficient use of the data. CV is widely used in spatial analyses; among many examples are Wang et al. (2022); Valavi et al. (2021); Kianian et al. (2021). However, the interpretation of the error estimates given by cross-validation is subtle even in the classical i.i.d. setting (Bates et al., 2023). We focus on the validation setting in this work due to its broad applicability and clear interpretation, though extensions to cross-validation, as well as a theoretical understanding of how the resulting error estimate relate to predictive performance is a promising direction for future work.
A.2 Covariate Shift
@@
In covariate shift, it is generally assumed that $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P^{\text{val}}$, $S^{\textup{{test}}}_{m}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}Q$ and $U^{\ell}_{i}|S^{\ell}_{i}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}\mu$ for $\ell\in\{\text{val},\text{test}\}$. Moreover, it is typically assumed the density ratio $\frac{dQ}{dP}$ exists and is bounded, although there is recent work relaxing this assumption in the context of non-parametric regression (Kpotufe and Martinet, 2021; Pathak et al., 2022). Mean estimation seeks to estimate $\mathbb{E}[U^{\text{test}}_{1}]$. Taking $(U^{\text{test}}_{m}|S^{\textup{{test}}}_{m})=\ell(Y^{\textup{{test}}}_{m},h^{\chi}(S^{\textup{{test}}}_{m}))$, this is the same task we consider, but with the (stronger) assumption that covariates are independent and identically distributed with a bounded density ratio between the test and validation distributions. This assumptions is not appropriate for validation with many spatial datasets, with particularly simple examples being when the task of interest requires prediction at a single spatial location, or on a regular grid. Our assumptions are essentially a conditional formulation of mean estimation under covariate shift. Many methods proposed for addressing mean estimation under covariate shift, including all the approaches we describe below, are based on re-weighting validation points, then applying the holdout approach described earlier with these weights.
@@
The kernel mean matching algorithm provides a solution under standard covariate shift assumptions with a bounded density ratio and assuming the average loss as a function of space $e_{h}$ lives in a reproducing kernel Hilbert space (RKHS) and has small norm (Gretton et al., 2009). Yu and Szepesvári (2012) provide a finite sample bound for this method
Appendix B ERROR ESTIMATION FOR MODEL SELECTION
@@
We are not able to formally characterize how SNN performs in model selection. But we provide rigorous results on rates of convergence of SNN that give suggestive guidance. In particular, (a) we first consider the case of a grid prediction task, where test sites are arranged on a grid. In this case, we are able to give a rate of convergence for both $1$NN and SNN (corollaries C.17 and C.18). If the validation data are i.i.d. our rate is faster than the minimax optimal convergence rate for predicting a Lipschitz function in the presence of additive, homoskedastic noise. Our result shows the same rate of convergence as Portier et al. (2024, Prop. 4), who considered test data that was i.i.d. instead of on a grid. Our result is suggestive that both nearest neighbor methods may perform well at model selection for validating maps (grid prediction).
(b) Second, we provide finite-sample bounds and asymptotic characterizations for general (non-grid) $Q^{\textup{{test}}}$ (corollary C.15). In this case, when validation data is i.i.d., up to logarithmic factors, we show that SNN converges at the optimal rate of convergence for Lipschitz functions (corollary C.19). Since in this case, the SNN convergence rate is not strictly faster than the training convergence rate, it might be difficult to select between statistical methods that converge at the optimal regression rate. But we would still expect to be able to select between (i) a statistical method that converges at the optimal rate and (ii) one that does not (for example a misspecified parametric model). Since the holdout and 1NN may not even be consistent, we could not rely on them to select the better model even in this latter, easier case.
Appendix C PROOFS OF CLAIMS
We now present results and proofs not included in the main text. We essentially follow the order of results in the main text. Section section C.1 gives sufficient conditions for assumption 2.5 to hold for a homoskedastic, additive noise model and squared loss for responses taking values in $[0,1]$, mentioned in section 2.2. Section C.2 focuses on the proof of proposition 3.3, first recalling several properties of covering numbers that will be used in the result. In section C.3 we restate and prove our propositions 4.1, 4.4 and 4.5, which show limitations of existing validation methods in the spatial setting we study. In section C.4 we prove a general result upper bounding the error in estimating the risk using $k$-nearest neighbors, as well as an upper bound on the error for $k^{\star}_{T_{2}}$. Section C.5 establishes the consistency of spatial nearest neighbors. Section C.6 discusses issues related to model selection and proves rates of convergence for spatial nearest neighbors.
C.1 Lipschitz Constant for Lipschitz Response and Predictive Method
@@
C.2 Proof that Independent and Identically Distributed Data Implies Infill Asymptotic with High Probability
@@
Our proof is similar to earlier proofs in Reznikov and Saff (2015, Section 5.2) and essentially follows the stackoverflow response Pinelis (2021) keeping track of numerical constants. Essentially, the idea is that were the fill distance to be large, there must be a ball of large radius that doesn’t contain any points in the sample of location. But because the probability of a point in the sample falling in any ball is bounded below in terms of the volume of the ball, this must be improbable as the sample size grows.
C.2.1 Preliminary Definitions and Lemmas Related to Nets and Covering Number
@@
Definition C.2 (Net, Covering Number).
@@
Lemma C.4 (Covering number of Unit Cube).
@@
By the upper bound in Wainwright (2019, Lemma 5.7),
@@ C.2.2 Main Proof
Returning to eq. 17 and taking a union over all elements in the net:
@@
where $B_{d}=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2}+1)}$ is the volume of the $d$-dimensional unit sphere. Equation 20 uses eq. 18. We now upper bound the terms in eq. 21.
Applying, propositions C.3 and C.4
@@
C.3 Proof of Inconsistency of Existing Methods
@@ C.3.1 Inconsistency of the Holdout
Proposition C.5 (Counterexample to the consistency of holdout).
@@
Combining eq. 27 and eq. 28, for all $N\geq N_{0}$ with probability at least $1/2$
@@
C.3.2 Inconsistency of $1$-Nearest Neighbor Estimator
We now turn to $1$-nearest neighbor risk estimation and restate and prove proposition 4.4: See 4.4 We again actually prove a strong result
Proposition C.7.
@@ C.3.2 Inconsistency of $1$-Nearest Neighbor Estimator
Take $t=\frac{A^{2}-C}{\delta}$ which is greater than 00 because $\delta\in(1-\frac{C}{A^{2}},1)$. Then eq. 33 becomes,
@@
Proof of Inconsistency of $1$-nearest neighbor
We return to our proof of proposition 4.4. The idea will be to consider a point prediction task and then apply proposition C.8 to show that with some fixed probability, the $1$-nearest neighbor estimator is a fixed distance away from the test risk, even as $N^{\textup{{val}}}$ increases.
@@
We now apply proposition C.8 with $U=|R_{Q^{\textup{{test}}}}(h)-(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)|$ to conclude that for $\delta\in(1-\frac{VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}}{\Delta^{2}},1)$ with probability at least $1-\delta$
@@
As neither $\delta$ nor the right hand side of eq. 43 depend on $N^{\textup{{val}}}$, $1$-nearest neighbor is not consistent under infill asymptotics. ∎
Proposition C.9.
Consider $\mathcal{Y}=\{0,1\}$ and $\ell(a,b)=\begin{cases}0&a=b\\ 1&\text{otherwise}\end{cases}$, $\mathcal{S}=[0,1]^{d}$ and $S^{\textup{{val}}}_{m}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}\mu$, where $\mu$ is any measure with density with respect to Lebesgue measure. Fix any $S\in\mathcal{S}$ and any predictive method $h$ and let $Q^{\textup{{test}}}=\delta_{S}$. Then,
@@
Because $S^{\textup{{val}}}_{m}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}\mu$, and $\mu$ has Lebesgue density, the nearest neighbor to $S$ is almost surely unique. This implies that $\hat{R}_{\textup{{NN}},1}(h)\in\{0,1\}$. Therefore,
@@
C.3.3 Inconsistency of Nearest Neighbors with Number of Neighbors Depending On Number of Validation Points